# Knowledge Prompting: How Knowledge Engineers Use Large Language Models

### v1

@@ 3. Methodology
During August $20232023$ we organised a four-day research hackathon for knowledge engineers and AI researchers to investigate KE with prompt engineering.333urlhttps://king-s-knowledge-graph-lab.github.io/knowledge-prompting-hackathon/ The hackathon format offered a tool to explore emerging practices as KE processes and approaches are disrupted by generative AI. This hackathon idea emerged from discussions at the Dagstuhl Seminar on “Knowledge Graphs and their Role in the Knowledge Engineering of the 21st Century” held in September 2022. The project topics (see Table 1) were defined in a community process with $3030$ experts (Groth et al., 2023). The hackathon hosted $3939$ participants from $1515$ different institutes and various environments from academia and industry, with $3131$ PhD students, $22$ postdocs, $11$ lecturers, $22$ professors, and $33$ industry members. The hackathon attendees were selected based on their background and experience with KE. They were selected from European research labs that have a strong profile in either publishing in KE venues, or in offering well-used KE industry products. During the hackathon, participants split into $77$ groups of $55$ to $77$ members, based on their prior experience in KE. Each group investigated one of the topics presented in Table 1. Participants received high level descriptions of the projects but were free to choose which tools, methodologies, domains, and LLM techniques they focused on.
@@
Table 1. Description of hackathon topics with associated tools and methodologies. Appendix C Table 7 includes descriptions and references for tools and methodologies included in this table.
Hackathon topic KE Tools/Methodologies LLM Tools/Methodologies
@@
Determine if LLMs can extract knowledge structures, including inference rules, to go with facts for KG construction Triples investigation ChatGPT, Few-shot prompt
@@
Create a framework providing tools for collaborative human-AI ontology engineering and requirements elicitation Competency Questions, eXtreme Design methodology ChatGPT, Multiple prompting techniques
@@
Determine how KE tasks can be supported from LLMs NeOn methodology, Competency Questions, HermiT Reasoner, OOPS PaLM, Llama, ChatGPT, Few-shot prompt
Determine if LLMs perform reasoning tasks completely in natural language Triples investigation ChatGPT, Multiple prompting techniques
@@
Determine if LLMs can perform ontology alignment (i.e. identify and match entities between ontologies) OAEI ChatGPT, Zero-shot prompt
@@
Determine if multimodal LLMs can be used towards the construction of multimodal KGs Investigate triples mPLUG-Owl, InstructBLIP, Text only prompt (no text + image)
@@
Determine if we can perform ontology refinement (i.e. techniques for KG completion and correction) using LLMs OntoClean ChatGPT, Llama, Claud, Few-shot prompt
3.1. Data Collection and Preparation
@@
Data was collected using three approaches to ensure rich data. The complete list of datasets can be seen in Table 2.
@@ 1. Ethnographic observation
Two researchers performed an ethnographic observation study, using the ‘observer as participant’ technique for ethnographic research (emerald publishing limited, 2023). The observer as participant role moves the observer into the study environment and closer to the activity of interest. The participants are aware of the observer, but the observer is not engaging with the participants and does not ask or interact with anyone. The researchers observed the various groups and kept detailed notes of the $77$ groups’ interactions and decisions.
@@ 2. Documents collected from the hackathon
Research output in the form of reports and documentation from the hackathon was collected. Hackathon groups produced reports regarding their experience answering specific questions and slide presentations with their strategies and results. This resulted in $77$ ppts and reports
@@ 3. Semi-structured interviews
After the hackathon, we contacted $1414$ participants who had volunteered to participate in follow-up interviews to talk about their opinions and experiences interacting with LLMs for KE tasks. Interviewees had backgrounds in KG construction, KG explainability, ontologies, and machine learning, as well as biology and bioinformatics, data management, FAIR data (Dodds, 2020), and digital humanities. We conducted these interviews virtually between 17/8/2023178202317/8/202317 / 8 / 2023 and 6/9/20236920236/9/20236 / 9 / 2023 using Microsoft Teams. [^4] https://www.microsoft.com/en-gb/microsoft-teams/group-chat-software The interview guide can be found in Appendix A.
We used Otter.ai  [^5] https://otter.ai/ to transcribe the interview recordings. This and the written data sources were then uploaded into NVIVO  [^6] https://help-nv.qsrinternational.com/20/win/Content/about-nvivo/about-nvivo.htm
Table 2. Number of documents and descriptions for each category of documents in the dataset.
Selected method # documents
Ethnographic study $22$ set of observation notes from two of the authors
Hackathon reports $77$ reports from the Hackathon groups
Hackathon presentations $77$ slide presentations from the Hackathon groups
Interviews $1414$ semi-structured interviews with Hackathon participants
3.2. Data Analysis
To address our research questions, we conducted thematic analysis across all the data. We first used an inductive method to search for key themes from the interview topic guide (see Appendix A). Table 3 shows the main themes identified. We then used a deductive method. We read the data corpus again for grounded themes that emerged in the data. The final codebook is shown in Appendix B.
@@ 3.2. Data Analysis
Table 3. Themes identified through the inductive analysis.
Themes Description
@@
Evaluation Evaluation issues and suggestion for LLM outputs
Skills The skills and qualities participants have or need to gain to interact with LLMs
Challenges Challenges that participants face while using LLMs fro KE tasks
@@
LLM interaction opinions Participants opinion about the use of LLMs for KE tasks
3.3. Ethics
The ethnographic study was approved by Authors’ institution’s Ethical Advisory Committee via the Full Application Form. Informed electronic (using Microsoft Forms [^7] https://forms.office.com/Pages/DesignPageV2.aspx and storing in CSV) consent was given by the participants. Thirty out of the $3939$ hackathon participants consented to the ethnographic study. Participants who did not consent were not observed. No personal information was collected in the analysis. The documentary data were collected as an output of the hackathon.
@@ 4. Results
4.1. What are the main challenges experienced by knowledge engineers when using LLMs for knowledge engineering tasks?
@@
Testing the use of LLMs for KE tasks was been challenging for all participants. Our observations during the hackathon revealed that the main concerns were identifing appropriate dataset to use, prompting LLMs efficiently, and evaluating the LLM outputs. We asked interviewees which of these they considered important.
@@
Dataset.Interviewees (1), (14), (12) emphasised the importance of datasets as a starting point, and groups used prompting as a means to create a dataset (Observer 1). Without a dataset, they felt they could not continue with prompting and evaluation (14). “There are many codes online…to interact with language models that wasn’t any problem at all. And the evaluation given that you know, you possess a ground truth,…was not too difficult. So, the most challenging case work for us was finding a dataset” (8). Finding a dataset was challenging for specific group tasks mainly because of the time pressure (14) and fast planning (1) needed to find and set domain knowledge for the task. On the other hand, an interviewee noted that LLMs should support establishing datasets. “The dataset is not what concerns me because if anything, I’m pretty impressed with what large language models potentially could do in terms of extracting, like I was saying, concepts, relationships, maybe even constraints or parameters or universal restrictions or existential restrictions, from unstructured data” (13).
@@
Prompting. An interviewee believed prompting is challenging due to some knowledge engineers’ lack of NLP training and experience, “I think prompting templates we’re going to be, generally more difficult for Semantic Web people, because we are not necessarily natural language processing people” (1). This was also emphasised in the skills that a knowledge engineer needs to work efficiently with LLMs. Furthermore, many participants noted the time-consuming phase of iteratively testing prompts to receive desirable outcomes (3), (Observer 1). Some suggested the usefulness of possible “templates for prompting” (Observer 1), (Observer 2), (3), (13). In addition, “once we settled on a specific prompt, we said okay, it seems like it works good enough, the most tricky part of it was to make the language model consistent in its responses” (3). This was particularly problematic when the process was automated with thousands of prompting iterations. Furthermore, syntactical errors also break the scripts and prevent automation (3). In contrast, an interviewee pointed out that prompting engineering is “less scientific” and models “with billions of parameters are not controllable,…, resulting in an output that you cannot reproduce one hour later, two days later. So, this might be a problem” (9).
@@
Evaluation. Participants believed that evaluation is challenging mostly due to the lack of benchmarks for specific KE activities, which depends on manual intervention, “so there are papers doing auto prompting…I’m a little biased towards evaluation simply because in my tasks evaluation pretty much boils down to a manual evaluation, there’s no other way to do it…it’s not that we have solved the dataset problem, but we can synthesise some datasets, or we can you know, create them artificially somehow. Evaluation still remains a challenge because. it has to be done manually” (5). In the case of using a specific controlled dataset for KG construction, the accuracy and coverage of the KG can be automatically evaluated (2), (10). However, in tasks such as designing competency questions or an ontology, there is not always a gold standard to compare with, and manual work from an expert is needed to evaluate the output. Nevertheless, exploring the use of holistic LLM benchmarks such as HELM (Bommasani et al., 2023) was suggested (5).
4.2. How do knowledge engineers evaluate LLMs’ output for their practices?
@@
We further asked interviewees whether current evaluation techniques for KE tasks could be used to evaluate LLM outputs. Most respondents believed that evaluation mainly depends on the task, but highlighted that current evaluation metrics are not sufficient for many KE activities. Following this, we asked them what a new benchmark or metric would look like.
@@
Current evaluation techniques. Some interviewees consider that for specific tasks, like using a controlled dataset (13),(14) and designing ontologies with simple taxonomies (12), evaluation metrics like F1 score, precision and recall (8), and comparison with gold standards are sufficient metrics for evaluation. However, others felt that regarding KE task evaluation, “it’s not [only] a problem that concerns only interaction with large language models, it’s a broader problem that needs to be solved” (9). Many KE tasks, like ontology semantic and reasoning errors, require human evaluation (1), (5) because “there is no automatic way [to assess]…how well the ontology represents the domain knowledge” (11), and this is important because “if it’s mistaking…in relations which creates the hierarchy…[that] is much more important than making a mistake on some data.” (2)
@@
In addition, others highlighted the need to evaluate the additional knowledge i.e., the extra knowledge produced by LLMS (10) and the knowledge is missing from KGs (4),(6),(14). An interviewee (10) highlighted the extra knowledge produced by LLMs based on their training, for example, when we asked the LLMs to create triples from an expression like “X was born in Y”, we expected “X-is a-person”, “Y-is a-country”, and “X-birthplace-Y, but we also may get “Y-part of-European Union”. These outputs include knowledge that may be needed to complete a KG, but it requires evaluation. In addition, a way to evaluate the knowledge missed in the LLMs’ outputs is needed. An interviewee pointed out that, “by using these classical evaluation metrics…[we check] what is inside the [input] text and [we miss] what is not in the text” (4).
@@
Finally, interviewees emphasised the difficulty in processing LLMs output mainly because LLMs return “strings” (10) of text and we lack a way to evaluate the “generative text” (7). It is common in KE tasks to use a specific format like JSON or the expression of an ontology. Even in the cases where participants asked the LLM to give output in these particular formats, they faced difficulties because, “the generated string contains inconsistencies, you really need to parse it and get what you want from the string itself, and then you need to convert it into a list, whatever you need” (10). Complicating this is the fact that, LLM outputs are not all the same type (4), for example, a KG includes dates, values, images, URLs etc. It is not possible to have a united way to evaluate the different types. Furthermore, another factor in output evaluation is the similarity of the results (4). Current evaluation techniques work on a match, not similarity, meaning they do not manage variations well, and LLMs are prone to inconsistency.
@@
New evaluation techniques. Most of the interviewees noted that current evaluation techniques are not sufficient and suggested possible alternatives. Some focused on the ontology design (2), (5), suggesting the development of a set of ontologies to be used as gold standards (observer 1). Others suggested the creation of toolkits (1) for LLMs to review ontology errors similar to existing ontology toolkits, like OntoClean (Guarino and Welty, 2009) and Oops (Poveda-Villalón et al., 2014). However, they acknowledged that this cannot be applied to large-scale KG like Wikidata, and manual work will still be required. Humans could be more efficient in evaluation if they have the information which prompt is related to the results (Observer 1).
@@
Another suggestion was to test the consistency of data and semantics in a real scenario, “for a company using this data and querying this data, and the simplicity of querying this data, and also the performances of these knowledge graphs in queries per second. And also the errors and the consistency of the data that is created,…making sure that the semantics of the data stored in the knowledge graph is consistent and reliable” (9).
@@
In addition, some interviewees were more specific, suggesting the use of techniques used in other fields, like fact-checking (10), adversarial algorithms (13), and self-play from reinforcement learning (12). An interviewee suggested using fact-checking techniques to establish that the extra knowledge received is correct and can be used for the KG construction (10). Another interviewee was fascinated by the concept of adversarial algorithms, a technique used to test machine learning by misguiding a model with malicious input. They suggested using this technique to test LLM results (13). In addition, another suggested the use of self-play technique, a reinforcement learning technique for improving agents’ performance by playing “against themselves”. They suggested to “let the model come up with, a whole host of answers, then compare them between each other [and]…get a more refined answer”, but this will require “more time and computing” (12).
@@
Two interviewees (3), (6) commented that we should see LLMs as assistants making the processes in KE smoother and faster because, “we know that as humans, we are lazy to wrote down some stuff like the competency question, so, if [you have a suggestion of] competency question you decide [faster], actually, which one you want to use” (6) and the same can apply every step of the way. They suggested that evaluation should focus on human satisfaction, “so the language model should probably optimise not really the performances, but the satisfaction of the user at least dealing with that” (3).
@@
4.3. What skills does a knowledge engineer need to incorporate LLMs into their practice?
@@
Developing a scientific framework was found fro one interviewee to help define the task’s workflow, describing it as “defining objective, defining hypotheses, defining an evaluation strategy, designing so an experiment” (7).
@@
On the other hand some interviewees were interested in gaining more experience to go beyond trial-and-error (4), (3), (1), (14). One interviewee acknowledged that prompting can be an intricate process. Some interviewees felt that time for experimentation with large language models is important to build an understanding of prompting and this familiarity with LLMs can improve performance (5). This can also allow for quicker execution of prompting to streamline the workflow, one interviewee noting, “I’m pretty sure that a lot of tricks or tips that I don’t know about” (6).
@@
4.4. How aware are knowledge engineers of bias as an LLM/KG safety issue?
@@
Particular issues related to LLMs were noted, for instance, ”there are so many confounding issues in the training of an LLM that even if you address bias in a prompt you can’t see the resources it is using to make a decision” (6). This respondent suggested that there could be ways to mitigate this by mapping the results of repeated prompting. One noted that harm can arise from use, even in a well-tested model. “So even though [LLMs] pass [bias] tests, I think there’s still risk if you prompt them in a hateful or discriminative way” (10).
@@
Some interviewees felt there might be a tension between attempting to reduce bias and model efficacy (2), (12) ”There are evidences that if you debias a language model, it usually downgrades in performance” (8). Whereas most respondents felt bias was found in data, a few perceived bias might also lie in both the LLM (7) and KGs, specifically in the semantics (9). One respondent felt that safety was not a concern in KGs (14).
4.5. What factors may affect knowledge engineers’ trust and uptake of the LLM technology?
Interviewees have varied opinions on the use of LLMs to their practices. Some find their use promising, and others are more sceptic.
Promising. Interviewees believe that LLMs cannot replace humans (3), but they can support KE activities. Using LLMs we can reduce time spent on tasks (11), (hackathon report), improve communication between experts in different disciplines (hackathon report), perform tasks even without extended experience (12), for example, in ontology design, and incorporate NLP pipelines in our practices without the need to develop them from the beginning (7). However, “we need to use them carefully by providing human oversight, [or] human in the loop, or our pipelines need to add options [where] human can intervene… to avoid passing those errors produced by large language models to the downstream tasks” (1).
@@ 4.5. What factors may affect knowledge engineers’ trust and uptake of the LLM technology?
Sceptic. Some interviewees were found to be more sceptical about the use of LLMs for KE tasks. They felt that LLMs by training are not up-to-date with the latest knowledge and that hallucination is very common (1). Moreover, LLMs respond with a prediction of what they think is the answer without giving information about how “confident” they are in the answer (13). In reverse, trusted KGs already support AI with trusted knowledge and can also support LLMs (1).

### v3

@@ 3 Methodology
During August 2023 we organised a four-day research hackathon for knowledge engineers and AI researchers to investigate KE with prompt engineering. [^4] https://king-s-knowledge-graph-lab.github.io/knowledge-prompting-hackathon/ The hackathon format offered a tool to explore emerging practices as KE processes and approaches are disrupted by generative AI. The project topics (see Table 1) were defined in a community process with 30 experts [58]. The hackathon hosted 39 participants from 15 different institutes and various environments from academia and industry, with 31 PhD students, 2 postdocs, 1 lecturers, 2 professors, and 3 industry members. The hackathon attendees were selected based on their background and experience with KE. They were from European research labs that have a strong profile in either publishing in KE venues, or in offering well-used KE industry products. During the hackathon, participants are divided into 7 groups of 5 to 7 members, based on their prior experience in KE. Each group explored one of the topics listed in Table 1. While participants were provided with high-level project descriptions, they had the flexibility to select the tools, methodologies, domains, and LLM techniques they wanted to focus on.
@@
Table 1: Overview of hackathon topics with associated tools and methodologies. Detailed descriptions and references for the tools and methodologies listed here can be found in Appendix C and Table 7.
Hackathon Topics KE task KE Tools & Methodologies GenAI Tools & Methodologies
@@
Determine if generative AI can extract knowledge structures, including inference rules, to go with facts for KG construction Use prompting to fill in triples, for example Coldplay - BandHasMember - ? Triples investigation ChatGPT, Few-shot prompt
@@
Create a framework providing tools for collaborative human-AI ontology engineering and requirements elicitation If there is a user story, use LLMs to create competency questions and then build an ontology Competency Questions, eXtreme Design methodology ChatGPT, Multiple prompting techniques
@@
Determine how KE tasks can be supported by generative AI Use ChatGPT to re-create the wine ontology NeOn methodology, Competency Questions, HermiT Reasoner, OOPS! PaLM, Llama, ChatGPT, Few-shot prompt
Determine if generative AI perform reasoning tasks completely in natural language Develop an agent (i.e. LLM) to collaboratively create a comprehensive knowledge graph with a domain expert. This involves crafting both the graph’s structure and content interactively, with the agent taking the lead Triples investigation ChatGPT, Multiple prompting techniques
@@
Determine if generative AI can perform ontology alignment (i.e. identify and match entities between ontologies) Given a request to an LLM, explore whether two items in WIkidata should be aligned OAEI ChatGPT, Zero-shot prompt
@@
Determine if generative AI can be used towards the construction of multimodal KGs Given a painting, explore whether LLMs can fill in a KG with information, such as painting information, painter’s image, painter’s information, art historian Investigate triples mPLUG-Owl, InstructBLIP, Text only prompt (no text + image)
@@
Determine if we can perform ontology refinement (i.e. techniques for KG completion and correction) using generative AI Use OntoClean within an iterative process to improve an ontology and to better represent knowledge OntoClean ChatGPT, Llama, Claud, Few-shot prompt
3.1 Data Collection and Preparation
@@
Data was collected using three approaches to ensure comprehensive coverage. A detailed list of datasets is provided in Table 2.
Table 2: Number of documents and descriptions for each category of documents in the dataset.
Selected Method
Collected Documents
Ethnographic study 2 sets of observation notes from two of the authors
Hackathon reports 7 reports from the Hackathon groups
Hackathon presentations 7 presentations and slides from the Hackathon groups
Interviews 14 semi-structured interviews with Hackathon participants
@@ 1. Ethnographic observation
Two researchers performed an ethnographic observation study, using the ‘observer as participant’ technique for ethnographic research [94]. The observer as participant role moves the observer into the study environment and closer to the activity of interest. The participants are aware of the observer, but the observer is not engaging with the participants and does not ask or interact with anyone. The researchers observed the various groups and kept detailed notes of the 7 groups’ interactions and decisions.
@@ 2. Documents collected from the hackathon
Research output in the form of reports and documentation from the hackathon was collected. Hackathon groups produced reports regarding their experience answering specific questions and slide presentations with their strategies and results. This resulted in 7 presentations and reports.
@@ 3. Semi-structured interviews
After the hackathon, we contacted 14 participants who had volunteered to participate in follow-up interviews to talk about their opinions and experiences interacting with generative AI for KE tasks. Interviewees had backgrounds in KG construction, KG explainability, ontologies, and machine learning, as well as biology and bioinformatics, data management, FAIR data [38], and digital humanities. We conducted these interviews virtually between 17 August 2023 and 6 September 2023 using Microsoft Teams. [^5] https://www.microsoft.com/en-gb/microsoft-teams/group-chat-software The interview guide can be found in Appendix A.
We used Otter.ai [^6] https://otter.ai/ to transcribe the interview recordings. The transcribed data sources were then uploaded into NVIVO [^7] https://help-nv.qsrinternational.com/20/win/Content/about-nvivo/about-nvivo.htm.
3.2 Data Analysis
To address our research questions, we conducted thematic analysis across all the data. We first used an inductive method to search for key themes from the interview topic guide (see Appendix A). Table 3 shows the main themes identified. We then used a deductive method. We read the data corpus for grounded themes that emerged in the data. The final codebook is shown in Appendix B.
@@ 3.2 Data Analysis
Table 3: Themes identified through the inductive analysis.
Themes
Description
@@
Evaluation Evaluation issues and suggestions for generative AI outputs
Skills The skills and qualities participants have or need to interact with generative AI
Challenges Challenges that participants face while using generative AI for KE tasks
@@
Generative AI interaction opinions Participants opinion about the use of generative AI for KE tasks
3.3 Ethics
The ethnographic study was approved by Authors’ institution’s Ethical Advisory Committee via the Full Application Form. Informed electronic (using Microsoft Forms [^8] https://forms.office.com/Pages/DesignPageV2.aspx and storing in CSV) consent was given by the participants. Thirty out of the 39 hackathon participants consented to the ethnographic study. Participants who did not consent were not observed. No personal information was collected in the analysis. The documentary data were collected as an output of the hackathon.
@@ 4 Results
4.1 RQ1: What are the main challenges experienced by knowledge engineers when using generative AI for KE tasks?
@@
Testing the use of generative AI for KE tasks was challenging for all participants. Our observations during the hackathon revealed that the main concerns for participants were identifying appropriate datasets to use, prompting LLMs efficiently, and evaluating the LLM outputs. We asked interviewees which of these they considered important.
@@
Dataset. Interviewees (1), (14), (12) emphasised the importance of datasets as a starting point, and groups used prompting as a means to create a dataset (Observer 1). A dataset was used by groups, for example, to design ontologies based on a domain, to extend existing KGs, and to test reasoning tasks and ontology alignment techniques. Without a dataset, they felt they could not continue with prompting and evaluation (14). “There are many codes online…to interact with language models that wasn’t any problem at all. And the evaluation given that you know, you possess a ground truth,…was not too difficult. So, the most challenging case work for us was finding a dataset” (8). Finding a dataset was challenging for specific group tasks mainly because of the time pressure (14) and fast planning (1) needed to find and set domain knowledge for the task. On the other hand, an interviewee noted that they felt LLMs would be supportive for establishing datasets. “The dataset is not what concerns me because if anything, I’m pretty impressed with what LLMs potentially could do in terms of extracting, like I was saying, concepts, relationships, maybe even constraints or parameters or universal restrictions or existential restrictions, from unstructured data” (13).
@@
Prompting. An interviewee believed prompting is challenging if knowledge engineers’ lack NLP training and experience, “I think prompting templates are going to be, generally more difficult for Semantic Web people, because we are not necessarily natural language processing people” (1). This was also emphasised in the skills that a knowledge engineer needs to work efficiently with LLMs. Furthermore, many participants noted the time-consuming phase of iteratively testing prompts to receive desirable outcomes (3), (Observer 1). Some suggested the usefulness of possible “templates for prompting”  [6, 193] (Observer 1), (Observer 2), (3), (13). In addition, “once we settled on a specific prompt, we said okay, it seems like it works good enough, the most tricky part of it was to make the language model consistent in its responses” (3). This was particularly problematic when the process was automated with thousands of prompting iterations. Furthermore, syntactical errors also break the scripts and prevent automation (3). In contrast, an interviewee pointed out that prompting engineering is “less scientific” and models “with billions of parameters are not controllable,…, resulting in an output that you cannot reproduce one hour later, two days later. So, this might be a problem” (9).
@@
Evaluation. Participants believed that evaluation is challenging mostly due to the lack of benchmarks for specific KE activities, which depends on manual intervention, “so there are papers doing auto prompting…I’m a little biased towards evaluation simply because in my tasks evaluation pretty much boils down to a manual evaluation, there’s no other way to do it…it’s not that we have solved the dataset problem, but we can synthesise some datasets, or we can you know, create them artificially somehow. Evaluation still remains a challenge because it has to be done manually” (5). In the case of using a specific controlled dataset for KG construction, the accuracy and coverage of the KG can be automatically evaluated (2), (10). However, in tasks such as designing competency questions or an ontology, there is not always a gold standard to compare with, and manual work from an expert is needed to evaluate the output. Nevertheless, exploring the use of holistic LLM benchmarks such as HELM [93] was suggested (5). Recent progress in developing benchmarks for LLMs in KE reflects the community’s strong emphasis on establishing standardised methods for their evaluation [131, 75].
4.2 RQ2. How do knowledge engineers evaluate generative AI output for their practices?
@@
We further asked interviewees whether current evaluation techniques for KE tasks could be used to evaluate generative AI outputs. Most respondents believed that evaluation mainly depends on the task, but highlighted that current evaluation metrics are not sufficient for many KE activities. Following this, we asked them what a new benchmark or metric would look like.
@@
Current evaluation techniques. Some interviewees consider that for specific tasks, like using a controlled dataset (13),(14) and designing ontologies with simple taxonomies (12), evaluation metrics like F1 score, precision and recall (8), and comparison with gold standards are sufficient metrics for evaluation. However, others felt that regarding KE task evaluation, “it’s not [only] a problem that concerns only interaction with LLMs, it’s a broader problem that needs to be solved” (9). Many KE tasks, like ontology semantic and reasoning errors, require human evaluation (1), (5) because “there is no automatic way [to assess]…how well the ontology represents the domain knowledge” (11), and this is important because “if it’s mistaking…in relations which creates the hierarchy…[that] is much more important than making a mistake on some data.” (2)
@@
In addition, others highlighted the need to evaluate the additional knowledge, specifically the extra knowledge produced by LLMS (10) and the knowledge is missing from KGs (4),(6),(14). An interviewee (10) highlighted the extra knowledge produced by LLMs based on their training, for example, when we asked the LLMs to create triples from an expression like “X was born in Y”, we expected “X-is a-person”, “Y-is a-country”, and “X-birthplace-Y, but we also may get “Y-part of-European Union”. These outputs include knowledge that may be needed to complete a KG, but it requires evaluation. In addition, a way to evaluate the knowledge missed in the LLMs’ outputs is needed. An interviewee pointed out that, “by using these classical evaluation metrics…[we check] what is inside the [input] text and [we miss] what is not in the text” (4).
@@
Finally, interviewees emphasised the difficulty in processing LLMs output mainly because LLMs return “strings” (10) of text and thry lacked a way to evaluate the “generative text” (7). It is common in KE tasks to use a specific format like JSON or the expression of an ontology. Even in the cases where participants asked the LLM to give output in these particular formats, they faced difficulties because, “the generated string contains inconsistencies, you really need to parse it and get what you want from the string itself, and then you need to convert it into a list, whatever you need” (10). Recent advances in LLMs’ structured generation capabilities have sought to address this challenge, as such requirements have emerged across domains. Approaches such as token-level constrained decoding [83, 178, 39] and prompt-level output structuring [98] can enable LLMs to produce outputs in specific, predefined formats. Complicating this is the fact that LLM outputs are not all the same type (4), for example, a KG includes dates, values, images, URLs etc. It is not possible to have a united way to evaluate the different types. Furthermore, another factor in output evaluation is the similarity of the results (4). Current evaluation techniques primarily rely on exact matching rather than similarity, making them less effective at handling variations and leaving LLMs prone to inconsistency.
@@
New evaluation techniques. Most of the interviewees noted that current evaluation techniques are not sufficient and suggested possible alternatives. Some focused on the ontology design (2), (5), suggesting the development of a set of ontologies to be used as gold standards (observer 1). Others suggested the creation of toolkits (1) for LLMs to review ontology errors similar to existing ontology toolkits, like OntoClean [60] and Oops [136]. However, they acknowledged that this cannot be applied to large-scale KG like Wikidata, and manual work will still be required. Humans could be more efficient in evaluation if they have the information which prompt is related to the results (Observer 1). Recent advances in KG evaluation demonstrate that integrating LLMs with human-in-the-loop approaches can be effective [170].
@@
Another suggestion was to test the consistency of data and semantics in a real scenario, “for a company using this data and querying this data, and the simplicity of querying this data, and also the performances of these KGs in queries per second. And also the errors and the consistency of the data that is created,…making sure that the semantics of the data stored in the KG is consistent and reliable” (9).
@@
In addition, some interviewees were more specific, suggesting the use of techniques used in other fields, like fact-checking (10), adversarial algorithms (13), and self-play from reinforcement learning (12). An interviewee suggested using fact-checking techniques to establish that the extra knowledge received is correct and can be used for the KG construction (10). Another interviewee was fascinated by the concept of adversarial algorithms, a technique used to test machine learning by misguiding a model with malicious input. They suggested using this technique to test LLM results (13). In addition, another suggested the use of self-play technique, a reinforcement learning technique for improving agents’ performance by playing “against themselves". They suggested to “let the model come up with, a whole host of answers, then compare them between each other [and]…get a more refined answer”, but this will require “more time and computing” (12).
@@
Two interviewees (3), (6) commented that we should see generative AI as assistants making the processes in KE smoother and faster because, “we know that as humans, we are lazy to write down some stuff like the competency question, so, if [you have a suggestion of] competency question you decide [faster], actually, which one you want to use” (6) and the same can apply every step of the way. They suggested that evaluation should focus on human satisfaction, “so the language model should probably optimise not really the performances, but the satisfaction of the user at least dealing with that” (3).
@@
4.3 RQ3. What skills does a knowledge engineer need to incorporate generative AI into their practice?
@@
Developing a scientific framework was found from one interviewee to help define the task’s workflow, describing it as “defining objective, defining hypotheses, defining an evaluation strategy, designing so an experiment” (7).
@@
On the other hand some interviewees were interested in gaining more experience to go beyond trial-and-error (4), (3), (1), (14). One interviewee acknowledged that prompting can be an intricate process. Some interviewees felt that time for experimentation with generative AI is important to build an understanding of prompting and this familiarity with LLMs can improve performance (5). This can also allow for quicker execution of prompting to streamline the workflow, one interviewee noting, “I’m pretty sure that a lot of tricks or tips that I don’t know about” (6).
@@
4.4 RQ3. How aware are knowledge engineers of using generative AI responsibly?
@@
Particular issues related to LLMs regarding provenance and source were noted, for instance, ”there are so many confounding issues in the training of an LLM that even if you address bias in a prompt you can’t see the resources it is using to make a decision” (6). This respondent suggested that there could be ways to mitigate this by mapping the results of repeated prompting. One noted that harm can arise from use, even in a well-tested model. “So even though [LLMs] pass [bias] tests, I think there’s still risk if you prompt them in a hateful or discriminative way” (10).
@@
Some interviewees felt there might be a tension between attempting to reduce bias and model efficacy (2), (12) ”There are evidences that if you debias a language model, it usually downgrades in performance” (8). Whereas most respondents felt bias was found in data, a few perceived bias might also lie in both the LLM (7) and KGs, specifically in the semantics (9). One respondent felt that safety was not a concern in KGs (14). Ever since we conducted the interviews, work has surveyed biases and fairness in LLMs [48], specifically building taxonomies of papers that propose metrics, datasets, and methods for LLM-based bias evaluation; but without considering LLMs at the task of KG construction. Since 2024, the Semantic Web community [130] has studied LLMs for KG construction tasks, e.g. ontology alignment [9, 54], class membership [7], and conversational ontology engineering [191]; although these works recognise the importance of bias and fairness in using LLMs for KG, none of them address them directly.
4.5 RQ5. What factors may affect knowledge engineers’ trust and uptake of the generative AI technology?
Interviewees have varied opinions on the use of generative AI in their practices. Some find their use promising, and others are more sceptical.
Promising. Interviewees believe that generative AI cannot replace humans (3), but they can support KE activities. Using LLMs we can reduce time spent on tasks (11), (hackathon report), improve communication between experts in different disciplines (hackathon report), perform tasks even without extended experience (12), for example, in ontology design, and incorporate NLP pipelines in our practices without the need to develop them from the beginning (7) [64, 26]. However, “we need to use them carefully by providing human oversight, [or] human in the loop, or our pipelines need to add options [where] human can intervene… to avoid passing those errors produced by LLMs to the downstream tasks” (1).
@@
Sceptic. Some interviewees were found to be more sceptical about the use of generative AI for KE tasks. They felt that LLMs without using any search or retrieval-agumented generation are not up-to-date with the latest knowledge and that hallucination is very common (1). Moreover, LLMs respond with a prediction of what they think is the answer without giving information about how “confident” they are in the answer (13). Recent work has sought to address this through approaches such as confidence elicitation [180, 52], self-assessment [142], chain-of-verification [36], and repeated sampling [23]. In reverse, trusted KGs already support AI with trusted knowledge and can also support LLMs (1).